# Efficient Approximation Algorithms for Fair Influence Maximization under Maximin Constraint

### v1

@@ Introduction
The reason the maximin objective $\Phi(S)$ does not satisfy submodularity is that a group can also be influenced by nodes belonging to other groups. Recall that $\rho(G,\mathcal{C})$ measures the connectivity strength between groups, where $\rho\rightarrow 0$ indicates weaker connectivity, and $\rho\rightarrow 1$ indicates stronger connectivity. When $\rho\rightarrow 0$, the likelihood that a group $c$ is influenced by nodes from other groups becomes negligible. In this case, AGM-GS is expected to perform well, approaching the approximation bound established in Theorem 2. Conversely, when $\rho\rightarrow 1$, across-group influence becomes more prevalent. This could result in significant overlaps among the seed sets $S^{c}\in\mathcal{S}$ for different $c\in\mathcal{C}$, making $k^{\prime}>>\lfloor k/m\rfloor$, which in turn enhances the approximation bound provided in Theorem 1.
Experiments
Datasets
@@
We conduct experiments on five real-world networks with ground-truth group structures. (1) The AVC network (Wilder et al. 2018) models an obesity prevention intervention in the Antelope Valley region of California. Each node in the network represents a real person associated with three demographic information. In our experiment, we choose the gender feature to divide the network into Male (51%) and Female (49%). (2) The labeled ENZ network (Rossi and Ahmed 2015) models the relations between enzymes, and (3) the labeled PRO network (Rossi and Ahmed 2015) models the relations between proteins. Both the ENZ network and the PRO network are divided into three groups. The group distributions for ENZ are 48.3%, 49.4%, and 2.3%, and for PRO are 48.7%, 48.1%, and 3.2%. (4) The UVM network (Traud, Mucha, and Porter 2012) is built from the Facebook data in UVM (University of Vermont), where every node is a member of UVM. The network is preprocessed by Lin et al. (Lin et al. 2023) who remove the nodes without a user profile. It is divided by the Role feature into Faculty (12%) and Student (88%), or divided by the Grade feature into Senior (40%) and Junior (60%). (5) Similarly, the UCSC dataset (Traud, Mucha, and Porter 2012) is built from the Facebook social network data in UCSC (University of California at Santa Cruz), where every node is a member of UCSC. The network is either divided by the role feature into Faculty (10%) and Student (90%) or divided by the Gender feature into Male (45%) and Female (55%). The detailed statistics of those networks are given in Table 1.
@@
Baselines
In experiments, we compare our proposed AGM-US and AGM-GS with (1) IMM (Tang, Shi, and Xiao 2015): a traditional IM method that focus on maximizing the influence spread, without considering fairness; (2) Myopic (Fish et al. 2019): a heuristic method that iteratively picks the node with the smallest probability of being influenced as a seed; and (3) Greedy: a naive greedy approach that iteratively picks the node providing the greatest gain in the maximin objective. Although simple, the Greedy usually serves as a natural baseline with favorable empirical performance, despite lacking strong theoretical guarantees due to the non-submodularity of the objective.
Experiment Results
AVC
@@
Since Myopic focuses on individual-aware fairness, it performs significantly worse than other baselines in terms of group-wise maximin utility. Besides, We consistently observe the performance ordering: AGM-GS >> AGM-US >> Greedy, indicating that AGM-GS empirically surpasses the theoretical guarantee given by AGM-US. In addition, IMM performs comparably to Greedy when the budget $k$ is small, but its performance keeps approaching that of AGM-GS as $k$ increases. This is primarily due to the balanced group structure of the AVC network, where maximizing the overall influence spread tends to yield a higher worst-off utility.
ENZ & PRO
@@
In contrast to the results on the AVC network, IMM performs poorly on both ENZ and PRO, even underperforming Myopic on the PRO network. The reason lies in that IMM focuses on maximizing global influence spread, which tends to overlook the minority group. Meanwhile, AGM-GS still consistently outperforms all other baselines, and notably surpasses Greedy on the PRO network.
UVM & UCSC
Both UVM and UCSC datasets feature two different group structures: one (based on ”Grade” or ”Gender”) is balanced, with an approximate $\%:5050\%$ split, while the other (based on ”Role”) is highly imbalanced, with an approximate $10\%$$90\%$ ratio. Corresponding results are presented in Figure 3, which includes four subfigures.
@@
Similar to the results on the AVC network, Myopic yields the weakest performance across all cases. For UVM (Grade) and UCSC (Gender), where the group structures are balanced, all methods perform comparably except for Myopic. In contrast, on UVM (Role) and UCSC (Role) with imbalanced group distributions, it can be observed that AGM-GS $\geq$ Greedy >> IMM >> AGM-US. Notably, IMM performs relatively well on these datasets due to the high network density, where both UVM and UCSC have an average degree exceeding 50, allowing the seeds of IMM to propagate effectively across different groups.
Price of Fairness
@@
In this subsection, we evaluate the Price of Fairness (PoF) for the four fairness-aware algorithms. PoF measures the trade-off between fairness and overall influence spread, quantifying how much influence is sacrificed to achieve fairness. Specifically, we use the influence spread achieved by IMM as the baseline and compute the PoF of any seed set $S$ as $PoF(S)=\frac{\sigma(S^{\text{IMM}})-\sigma(S)}{\sigma(S^{\text{IMM}})}$ with $S^{\text{IMM}}$ denoting the seed set of IMM. The PoF results across different networks are summarized in Table 2, where a negative value indicates that the influence spread of that seed set even exceeds that of IMM.
Table 2: Price of Fairness (PoF). Bold indicates the best (lowest) value, and underlined italic denotes the runner-up.
Network Method $k=10$ $k=30$ $k=50$
AVC Myopic 55.29% 49.78% 41.60%
Greedy 14.66% 10.91% 9.83%
AGM-US 2.15% 1.78% 2.62%
AGM-GS 1.90% 1.84% 1.90%
UVM (Role) Myopic 66.13% 70.29% 70.82%
Greedy 2.49% 2.88% 2.73%
AGM-US 1.73% 3.54% 7.37%
AGM-GS 1.90% 0.66% 0.24%
UVM (Grade) Myopic 65.50% 70.72% 71.19%
Greedy 0.75% 1.63% 0.41%
AGM-US 0.69% -0.12% 1.60%
AGM-GS 0.39% 0.65% 0.99%
UCSC (Role) Myopic 76.92% 83.22% 82.75%
Greedy 6.42% 1.97% 1.63%
AGM-US 3.06% 6.59% 9.83%
AGM-GS -0.46% 1.19% 1.78%
UCSC (Gender) Myopic 77.26% 82.56% 83.83%
Greedy -0.10% -0.21% -0.23%
AGM-US -0.22% 0.04% 1.91%
AGM-GS 0.02% 0.21% 0.02%
ENZ Myopic 50.89% 46.27% 43.29%
Greedy 23.78% 14.72% 13.58%
AGM-US 19.77% 16.32% 14.00%
AGM-GS 15.80% 13.06% 11.38%
PRO Myopic 52.39% 57.00% 53.77%
Greedy 35.16% 28.39% 22.12%
AGM-US 24.93% 20.39% 18.13%
AGM-GS 23.88% 18.35% 14.90%
@@
Overall, our proposed AGM-GS consistently achieves the best performance, incurring the lowest influence loss while attaining higher fairness. The only exception occurs on the UCSC (Gender), where AGM-GS slightly underperforms Greedy and AGM-US. This is likely because the group connectivity $\rho$ in UCSC (Gender) is the highest among all datasets. As previously discussed, AGM-GS is more effective when $\rho\rightarrow 0$. Hence, its performance slightly deteriorates in scenarios with strong group connectivity.
Conclusion
@@
This paper studies the Fair Influence Maximization (FIM) problem under the maximin constraint, where the goal is to maximize the utility of the worst-off group in a social network. To address the challenge that the maximin objective is not submodular, we propose a novel two-step optimization framework, leveraging the fact that influence spread within any individual group remains submodular, as we formally prove. In the first step, we design Inner-group Maximization (IGM), which fully exploits inner-group submodularity to select high-quality seed nodes for each group with theoretical guarantees. In the second step, we develop Across-group Maximization (AGM), which coordinates the final seed selection based on the candidate lists produced by IGM. AGM introduces two strategies: AGM-US (Uniform Selection), which selects seeds in a column-wise manner across groups and achieves a $\frac{1}{m}(1-1/e-\varepsilon)$ approximation that is independent of group structure; and AGM-GS (Greedy Selection), which greedily selects the node with the highest marginal gain and achieves a $(1-1/e-\varepsilon)$ approximation when groups are completely disconnected. Experimental results on five real-world networks validate the effectiveness of our framework, where the proposed algorithms consistently outperform the naive Greedy baseline. Notably, AGM-GS usually achieves better empirical performance compared to AGM-US, providing stronger practical results than the theoretical approximation guarantee suggests.

### v2

@@ 3.3 Discussion on AGM and $\rho$
The reason the maximin objective $\Phi(S)$ does not satisfy submodularity is that a group can also be influenced by nodes belonging to other groups. Recall that $\rho(G,\mathcal{C})$ measures the connectivity strength between groups, with $\rho\to 0$ indicating weaker connectivity and $\rho\to 1$ stronger connectivity. When $\rho\to 0$, the likelihood that a group $c$ is influenced by nodes from other groups becomes negligible. In this case, AGM-GS is expected to perform well, approaching the approximation bound established in Theorem 2. Conversely, when $\rho\rightarrow 1$, across-group influence becomes more prevalent. This could result in significant overlaps among the seed sets $S^{c}\in\mathcal{S}$ for different $c\in\mathcal{C}$, making $k^{\prime}>>\lfloor k/m\rfloor$, which in turn enhances the approximation bound of AGM-US in Theorem 1.
4 Experiments
4.1 Datasets
We conduct experiments on seven real-world networks with ground-truth group structures.
@@ 4.1 Datasets
(1) The AVC network (Wilder et al., 2018) models an obesity prevention intervention in the Antelope Valley region of California. Each node in the network represents a real person associated with three demographic information. In our experiment, we chose the gender feature to divide the network into Male (51%) and Female (49%). (2) The labeled ENZ network (Rossi and Ahmed, 2015) models the relations between enzymes, and (3) the labeled PRO network (Rossi and Ahmed, 2015) models the relations between proteins. Both the ENZ network and the PRO network are divided into three groups. The group distributions for ENZ are 48.3%, 49.4%, and 2.3%, and for PRO are 48.7%, 48.1%, and 3.2%. (4) The UVM network (Traud et al., 2012) was built from the Facebook data in UVM (University of Vermont), where every node is a member of UVM. The network was preprocessed by Lin et al. (2023), who remove the nodes without a user profile. It is divided by the Role feature into Faculty (12%) and Student (88%), or divided by the Grade feature into Senior (40%) and Junior (60%). (5) Similarly, the UCSC dataset (Traud et al., 2012) was built from the Facebook social network data in UCSC (University of California at Santa Cruz), where every node is a member of UCSC. The network is either divided by the role feature into Faculty (10%) and Student (90%) or divided by the Gender feature into Male (45%) and Female (55%). (6) The GOW network (Cho et al., 2011) was collected from the public API of Gowalla, a location-based social networking service in which users share their locations via check-ins. Nodes are classified based on their check-in records into check-in users (54.5%) and non-check-in users (45.5%). (7) The TWI network (Sarkar and Rózemberczki, 2021) network was collected from the public API, where nodes represent Twitch users and edges denote mutual follower relationships. We partition the network based on the age feature into mature users (47.0%) and immature users (53.0%).
Detailed statistics of those networks are given in Table 1.
@@
GOW 196.6k 950.3k 0.4187 2 Check-info
TWI 168.1k 6797.6k 0.4547 2 Age
4.2 Baselines
In experiments, we compare our proposed AGM-US and AGM-GS with (1) IMM (Tang et al., 2015): a traditional IM method that focuses on maximizing the influence spread without considering fairness; (2) Greedy (Fish et al., 2019): a naive greedy approach that iteratively picks the node providing the maximum gain in the maximin objective; and (3) LP-Greedy (Spaeh and Miyauchi, 2025): an LP-based method that formulates the problem as multiobjective submodular maximization and solves it with Gurobi optimizer (Gurobi Optimization LLC., 2025), has been shown to outperform previous approaches (Udwani, 2018; Tsang et al., 2019).
It is worth noting that the primary time cost of our method arises from IMM. Our additional steps are lightweight: AGM-US runs in $O(k)$, while AGM-GS runs in $O(mk\theta)$.
4.3 Maximin Objective
AVC.
@@ 4.3 Maximin Objective
We consistently observe the performance ordering: AGM-GS >> LP-Greedy >> AGM-US >> Greedy, indicating that AGM-GS empirically surpasses the theoretical guarantee given by AGM-US. In addition, IMM performs comparably to Greedy when $k$ is small, but its performance keeps approaching that of AGM-GS as $k$ increases. This results from the balanced group structure, where maximizing the overall influence spread benefits to the worst-off group.
ENZ & PRO.
@@
In contrast to the results on the AVC network, IMM performs poorly on both ENZ and PRO. The reason lies in the fact that IMM focuses on maximizing global influence spread, which tends to overlook the minority group. Meanwhile, AGM-GS still consistently outperforms all other baselines, and notably surpasses LP-Greedy on the PRO network.
@@
UVM & UCSC.
Both UVM and UCSC datasets feature two different group structures: one (based on ”Grade” or ”Gender”) is balanced, with an approximate $\%:5050\%$ split, while the other (based on ”Role”) is highly imbalanced, with an approximate $10\%$$90\%$ ratio. Corresponding results are presented in Figure 3, which includes four subfigures. We attempted to run LP-Greedy on both UVM and UCSC. However, it failed to produce results after more than three days of computation, even with $k=10$.
@@
For UVM (Grade) and UCSC (Gender), where the group structures are balanced, all methods perform comparably. In contrast, on UVM (Role) and UCSC (Role) with imbalanced group structures, it shows that AGM-GS $\geq$ Greedy >> IMM >> AGM-US. Notably, IMM performs relatively well on these datasets due to the high network density, allowing the seeds of IMM to propagate effectively across different groups.
GOW & TWI.
GOW and TWI are two large-scale networks, where TWI is even denser than UVM and UCSC. For LP-Greedy, we obtained results only on GOW with $k=10$, which cost 28 hours of computation.
(a) GOW
(b) TWI
Figure 4: Comparison results on the GOW and TWI network.
Results in Figure 4 show that IMM, Greedy, and AGM-GS achieve comparable performance, with IMM being slightly worse. This phenomenon is consistent with the results on UVM (Grade) and UCSC (Gender), where the group structures are relatively balanced (approximately $45\%$$55\%$ for both GOW and TWI). Notably, although Greedy attains similar performance, it requires 14 hours to produce results on TWI, whereas AGM-GS requires only approximately 1.1 times the runtime of IMM, completing in tens of minutes.
4.4 Price of Fairness and Runtime
@@
Price of Fairness (PoF) measures the trade-off between fairness and overall influence spread, quantifying how much influence is sacrificed to achieve fairness. Specifically, we use the influence spread achieved by IMM as the baseline and compute the PoF of any seed set $S$ as $PoF(S)=\frac{\sigma(S^{\text{IMM}})-\sigma(S)}{\sigma(S^{\text{IMM}})}$ with $S^{\text{IMM}}$ denoting the seed set of IMM. The PoF results across different networks are summarized in Table 2, which also lists the runtime results for $k=50$.
@@
Generally, our proposed AGM-GS consistently achieves the best performance, incurring the lowest influence loss while attaining higher fairness. On UCSC (Gender), GOW, and TWI, the PoF of AGM-GS is comparable to Greedy, likely due to their higher group connectivity $\rho$. As discussed earlier, AGM-GS is most effective as $\rho\rightarrow 0$; Hence, its performance slightly deteriorates in scenarios with higher $\rho$.
Moreover, both AGM-US and AGM-GS exhibit excellent scalability, with average runtimes less than one-tenth of that of Greedy. In contrast, LP-Greedy incurs substantially higher computational cost than Greedy and fails to produce results on dense or large-scale networks.
Table 2: PoF for $k=10,30,50$ and runtime (s) for $k=50$. Bold indicates the best PoF, and underlined italic denotes the runner-up.
Network Method $k=10$ $k=30$ $k=50$ runtime
AVC Greedy 13.20% 6.26% 9.83% 4.1
LP-Greedy 2.14% 1.65% 2.93% 4285.8
AGM-US 4.11% 2.43% 2.95% 0.5
AGM-GS 4.24% 2.07% 2.38% 1.3
ENZ Greedy 23.78% 14.72% 13.58% 282.7
LP-Greedy 17.26% 13.64% 12.64% 1711.9
AGM-US 19.77% 16.42% 14.00% 29.3
AGM-GS 14.42% 11.59% 11.38% 78.5
PRO Greedy 35.16% 28.39% 22.12% 616.7
LP-Greedy 25.77% 18.91% 16.67% 2355.4
AGM-US 24.93% 20.39% 16.54% 43.9
AGM-GS 21.88% 18.00% 14.68% 114.1
UVM (Role) Greedy 2.49% 2.99% 2.78% 1008.6
AGM-US 1.73% 3.65% 7.41% 69.1
AGM-GS 1.90% 0.78% 0.29% 75.0
UVM (Grade) Greedy 0.54% 1.72% 1.82% 1098.6
AGM-US 0.49% 0.00% 1.89% 69.8
AGM-GS 0.18% 0.74% 1.28% 76.3
UCSC (Role) Greedy 6.52% 1.71% 1.57% 1269.5
AGM-US 3.17% 6.34% 9.77% 68.5
AGM-GS -0.35% 0.93% 1.72% 74.3
UCSC (Gender) Greedy 0.13% 0.14% 0.00% 1203.3
AGM-US 0.01% 1.99% 2.14% 102.8
AGM-GS 0.25% 0.21% 0.26% 112.7
GOW Greedy 0.98% 0.83% 1.24% 2920.5
LP-Greedy 1.39% - - -
AGM-US 7.00% 3.93% 4.36% 157.1
AGM-GS 1.15% 0.76% 1.19% 172.2
TWI Greedy 0.19% 0.22% 0.29% 49905.3
AGM-US 4.70% 3.73% 3.56% 2663.2
AGM-GS 0.13% 0.33% 0.39% 2682.1
5 Conclusion
@@
This paper studies the Fair Influence Maximization (FIM) problem under the maximin objective, which aims to maximize the utility of the worst-off group. To address the non-submodularity of the maximin objective, we propose a two-step optimization framework that leverages the submodularity of influence spread within individual groups, as formally established. The first step, Inner-group Maximization (IGM), generates high-quality group-wise seed candidates with theoretical guarantees, while the second step, Across-group Maximization (AGM), coordinates seed selection across groups. AGM also includes two variants: AGM-US, which provides a $\frac{1}{m}(1-1/e-\varepsilon)$ approximation independent of group structure, and AGM-GS, which achieves a $(1-1/e-\varepsilon)$ approximation when groups are disconnected. Extensive experiments on seven real-world networks demonstrate that our methods consistently outperform state-of-the-art baselines, with AGM-GS typically achieving superior empirical performance.